### v3

# Network Structure Trumps Neuron Dynamics: Insights from Drosophila Connectome Simulations

Abstract
Despite the success of artificial neural networks, the necessity of real network structures in simulating intelligence remains unclear. Utilizing the largest adult Drosophila connectome data set, we constructed a large-scale network communication model framework based on simple neuronal activation mechanisms to simulate the activation behavior observed in the connectome. The results demonstrate that even with simple propagation rules, models based on real neural network structures can generate activation patterns similar to those in the actual brain. Importantly, we found that using different neuronal dynamics models, all produced similar activation patterns. This consistency across different models emphasizes the crucial role of network topology in neural information processing, challenging views that rely solely on neuron count or complex individual neuron dynamics. Moreover, we test the influence of network reconnect rate and find that at least 1%’s reconnect rate will ruin the activation patterns appeared before. By comparing network distances and spatial distances, we found that network distance better explains the information propagation patterns between neurons, highlighting the importance of topological structure in neural information processing. To facilitate these studies, we developed real-time 3D large spatial network visualization software, bridging a crucial gap in existing tools. Our findings underscore the importance of network structure in neural activation and provide new insights into the fundamental principles governing brain functionality.
Keywords: brain network, network communication model, drosophila connectome, activation pattern
1 Introduction
The neural systems of intelligent beings maintain constant communication[1, 2]. The intricate mechanism of this communication unfolds as follows: specific neurons, often input neurons, receive stimulus signals from the external environment or other parts of the body, changing from a resting to an activated state. Subsequently, they relay these stimulus signals to the next tier of neurons through synapses. This process gives rise to macroscopic emergence across the entire neural system. Such communication fosters the emergence of intelligence in all intelligent life forms, including human beings. Consequently, comprehending and replicating this process has become an indispensable avenue for unraveling biological intelligence and realizing artificial intelligence[3].
Neural systems are interconnected networks[4], with neurons engaging in constant communication via synapses within the network’s topology. Some scientists think the brain’s unique and complex network structure is pivotal to the emergence of intelligence[5]. Many scientists posit that network science is crucial for deciphering the mechanisms underlying the flexible regulation of neural communication[6, 7, 8]. Determining how to leverage network science methodologies stands as a paramount challenge in contemporary neuroscience.
However, whether this network structure determines the emergence of intelligence remains a question. In recent years, artificial neural networks (ANNs) have taken rapid development in enormous fields[9, 10, 11]. A notable recent achievement is large language model (LLM). However, the structure of any ANNs does not follow the actual brain network structure strictly, which poses a challenge to the viewpoint that network structure inherently contains intelligence. Simultaneously, many scientists in machine learning adhere to the scaling law: the more parameters the network has, the greater the degree of emergent intelligence[12]. This viewpoint suggests that intelligence is solely related to the number of parameters a model has, rather than the unique network structure between neurons. Under this perspective, the reason biological brains generate intelligence is due to the number of neurons, and the unique network structure is a coincidence of evolution.
To demonstrate the importance of brain network structure, network scientists are constantly exploring and experimenting. An approach is brain network communication model[13]. Scientists wants to use brain network communication model to understand and describe the interaction and communication between different parts of the brain. This model aids in studying the structural and functional aspects of neuronal networks and how they coordinate to accomplish complex cognitive processes[14].
The first challenge of this approach is obtaining a complete and accurate mapping of the brain’s structure. For example, current imaging technologies, such as MRI, fMRI, and electron microscopy, have limitations in resolution and speed. While MRI and fMRI provide non-invasive imaging, their spatial resolution is not sufficient to map individual neurons and synapses. On the other hand, techniques like electron microscopy, which can achieve higher resolution, are extremely slow and generate enormous amounts of data that are challenging to process[15]. Modeling from the micro scale is necessary because modeling at the level of brain regions or nerve fibers does not include the most basic structural information and lacks the most complete physical foundation.
Moreover, current models often rely on prevalent metrics such as signaling cost and activation time to assess model quality[16]. Yet, these metrics fall short of verifying whether a network model accurately reflects the brain’s actual activation patterns, which is the fundamental goal of brain simulation. There is a need for evaluation criteria that can affirm whether network communication models genuinely mirror the brain’s communication processes.
Based on facts and challenges above, benefiting from a largest-to-date adult Drosophila connectome data[17], this study attempts to investigate how the type of structure of neurons could affect the activation patterns of a brain. Our approach involves crafting a large scale network communication model framework based on simple mechanisms, designed to simulate the activation behavior observed in the connectome. Specifically, we establish the network representation of the connectome. Then we design a large scale threshold propagation model of nerve cells: when a neuron received enough stimuli, it activates and subsequently transmits signals to other neurons. After a few iterations, the whole network emerges an activation pattern. We use threshold model and sigmoid model as the basic neuron model. We also design an improved LIF model as the basic neuron model for comparison. Additionally, we propose an evaluation criteria based on the ratio of actually activated neurons to the expected number of activation. Experiments reveals that, from threshold , sigmoid, and LIF neuron model, no matter activation mechanism is, our model has significant activation in the type of neurons that should be affected by the stimulus and slight activation in irrelevant neurons. Moreover, to supply more evidence of the vitalness of network structure, we disturb the network by the way of degree-preserving rewiring and and find the disturbed network can’t generate any correct activation pattern.
In addition, we test the sensitivity of disturb degree. For details, we record the activation rate of areas under different reconnect rates and find that 1%’s random change of network structure will cause a complete change in function.
To prove the network structure, rather than the spatial structure, matters in activation pattern generation, we calculate the average network and spatial distance of input and activation neurons and results reveal that the average network distance between input neurons and their corresponding activation areas is shorter than unrelated areas, which is a strong evidence of network matters in message propagation in brain.
Furthermore, when providing unilateral stimuli to input neurons, we observe a bilateral response in the Drosophila connectome, aligning with reality. Then we calculate the statistics properties and find that the left and right hemispheres have extremely similar degree, clustering coefficient, and eigenvector centrality vector coefficient. The vectors’ Pearson coefficient are up to 0.9986, 0.9988, and 0.9991, respectively.
We also developed a 3D network visualization software see in appendix F, utilizing HTML5’s Canvas and Three.js based on WebGL technology for large spatial network visualization effects. This tool enables real-time observation of each neuron’s communication process.
2 Results
First, we report the basic statistics of Drosophila connectome. After that, to test the performance of our model in terms of realism on the data set and determine whether different artificial mechanisms can achieve similar activation patterns when the network structure changes, we give different stimuli and record the activation pattern of different neurons. In addition, we give a unilateral stimulus (stimulus one hemisphere) and record the bilateral response of neurons. Besides, to add more neurons’ real dynamic details, we also add an improved LIF model as the basic neuron dynamic model and record the activation patterns under stimuli. Finally, to prove the network structure matters in brain network communication, we calculate the average physical and network distance between different areas.
2.1 Experimental Settings
The basic methodology of the experiment is described as follows. In this study, a continuous stimulus signal (binary 0-1 signal) is applied to a subset of input neurons (such as visual or olfactory neurons) in the brain. Subsequently, we compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons. The evaluation criterion is determined by calculating the ratio of neurons that ought to be activated to those that are actually activated in a specific area (see as Equ.26 in appendix B). Given that the brain functions as a type of shallow neural network[53], we document the outcomes at an iteration step size of 5. Moreover, as the input remains constant and no randomization parameters have been introduced, the outcomes of each experiment are consistent, eliminating the need for repetition. For messages in related neurons, please read appendix C.
In linear threshold model, we set $\sigma_{0}=0.8$. In the simulation of the LIF model, numerous hyperparameters need to be used. All the hyperparameters used are listed in the following table. Biological hyperparameters are set according to real biologic experiment test[54]. Current proportion const is the optimal parameter obtained under equidistant search.
Table 1: List of Hyperparameters
Hyperparameter Description Value
$\tau$ Membrane time constant 15.0 ms
$V_{\text{rest}}$ Resting membrane potential -65 mV
$V_{\text{th}}$ Spike threshold -45 mV
[…]
$dt$ time step 0.55 ms
$\alpha$ Current proportion const 0.03/0.045 (visual/olfactory)
2.2 Visualization and Statistic Properties of Drosophila Connectome
In this section we first present the visualization of Drosophila Connectome (Fig.1(a)). Other visualization cases and methods of visualization are presented in appendix F. Visually, the shape of the brain appears to be physically similar, but whether the network structures are similar remains unknown. So we further investigate the network statistics of two hemispheres. We calculated the degree, clustering coefficient, and eigenvector centrality distribution for every node in each hemisphere (Fig.1(d), (e) and (f)). The Pearson correlation coefficient of degree, clustering coefficient, and eigenvector centrality vector in each hemisphere is 0.9986, 0.9988, and 0.9991, respectively. According to Fig.1(d), (e) and (f), these statistical indicators for both hemispheres are extremely similar, indicating that the brain exhibits strong symmetry in network structure.
From Fig.1(c) we can see that the degree, i.e., the number of connections to the nodes in the network, exhibits a long-tail distribution, implying that while a large number of nodes possess few connections, a small number of nodes contains a large number of connections[37]. It implies the real neural network have some same statistic properties as social, electrical and transportation networks. Also, we plot the correlation curves between degree and cell surface area, volume, and cell length (Fig.1). The results indicate that the degree distribution is logarithmically linearly correlated with these characteristics. This logarithmic relationship reveals a fundamental principle in the organization of neural networks in the brain.
To elaborate, as the surface area, volume, or length of a neuron increases, the number of synapses (represented by the degree) it forms grows logarithmically rather than linearly. This non-linear scaling is crucial for maintaining efficient information processing in the brain while preventing an explosion in connectivity that would be computationally and energetically costly.
These characteristics of the brain can explain Weber-Fechner’s law[47]: for any sensory pattern, the perceived intensity is a logarithmic function of the physical intensity. Specifically, the size, length, and area changes can only cause a logarithmic change in the number of synapses the neuron has. This means that as a stimulus becomes stronger, the increase in neural activity or connectivity is proportionally smaller, mirroring the logarithmic perception of intensity described by Weber-Fechner’s law.
(a) Visualization of Drosophila Connectome
(b) Correlation between degree distribution and neuron characteristics
Figure 1: Visualization and statistic properties of Drosophila connectome data set
These statistic results are a valuable supplement to previous findings[48] and are crucial for for improved understanding of the Drosophila brain connectcome structure.
2.3 Activation Patterns under Different Type of Stimuli
Primarily, to validate the efficacy of our model in signal transmission between input and output neurons, we showcase the activation patterns of intermediate and output neurons throughout the information transmission process. This demonstration serves to substantiate that our model accurately reflects real-world scenarios, shaping a quasi-real activation pattern.
The specific idea for the experiment is following: we stimulate different types of neurons (such as visual input neurons, olfactory neurons, etc.), and record the rate that neurons that are activated in different brain areas. If a specific stimulus causes a significantly greater response than the average in the brain area that it is theoretically supposed to activate, then it is considered that the model can generate a quasi-real activation pattern in the brain. We give 2 types of input neurons (visual, olfactory) stimulus and record the activation patterns of 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory). When we give neuron a stimulus, we set the state of neuron as 1.
To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks. Specifically, we perform degree-preserving rewiring on a certain proportion of edges in the brain, which means perturbing the network structure while ensuring that the degree of each node remains unchanged. In Fig.2 we set reconnect rate $p=10\%$. For its sensitivity analysis, please read appendix S4.
Fig.2 shows the activation patterns of different models. From Fig. 2(a), (e), and (i), it shows visual stimulus cause a strong impact (significantly more than average) on visual neurons (optic, visual projection) while causing a slight impact (significantly less than average) on other type of neurons under all three kinds of models. Specifically, under visual stimulus, in response to visual stimuli, the optic (vision related area) exhibit a 36.39%, 4.69% and 9.69% ’s activate rates, respectively. At the mean time, no olfactory-related neurons are activated.
In addition, olfactory stimulus cause a strong impact (significantly bigger than average level) on olfactory neurons (ALLN, ALIN, ALPN, ALON), reaching the activation rate by 94.07%, 78.01%, and 69.75%, respectively. And it cause a slight impact (significantly less than average) on other type of neurons. For example, it only activates 0.04% of optic and 0.79% of visual_projection in threshold model (Fig.2(c)).
Although the activation rates of the three models in specific areas differ (due to different activation mechanisms), the significant difference in the activation ratio between correlated and uncorrelated regions indicates the emergence of similar quasi-real activation patterns across all models. This suggests that different activation mechanisms are not the fundamental reason for changes in network functionality.
As for the disturbing models, in Fig.2 (b), (d), (f), (h), (j) and (l), the quasi-real activation patterns totally disappear. Across all activation modes and all types of stimuli, no significant differences in activation regions were observed. For example, in the threshold model, the maximum difference in the activation ratios of all areas under visual stimuli was only 3.162%. Moreover, the range of the regional activation ratios versus the average activation ratio was only 6.32%. Additionally, in the sigmoid and LIF models, no significant activation was observed in any area, indicating that stimuli under network disturbance did not lead to any significant response in any area.
These results indicate that no matter neuron model is, it will lead to similar activation patterns which correspond to reality. After disturbing the network structure, the original quasi-real activation patterns in all models completely disappear. It means the real network structure matters than activation mechanisms in generating the quasi-real activation patterns. Different activation mechanisms are not the primary cause of this realistic activation mechanism; it is the brain’s unique network structure that plays the crucial role. This aligns with the studies of complex systems: individual capabilities are not the determining factors in complex systems; what matters is the organization and interaction structure among individuals. This perspective highlights the importance of individual relationships and the overall system structure in determining emergent phenomena.
(a) threhold, visual
(b) threhold, visual (after disturb)
(c) threhold, olfactory
(d) threhold, olfactory (after disturb)
(e) sigmoid, visual
(f) sigmoid, visual (after disturb)
(g) sigmoid, olfactory
(h) sigmoid, olfactory (after disturb)
(i) LIF, visual
(j) LIF, visual (after disturb)
(k) LIF, olfactory
(l) LIF, olfactory (after disturb)
Figure 2: Experiment of intrinsic neurons’ activation. For disturbing model, we conduct 5 repeated experiments and take the average and standard deviation.
2.4 Distances between Input and Activation Areas
To further investigate how the network structure, rather than the spatial distribution of neurons, or the parameters of the network model or activation mechanisms, predominantly determines the dynamics of brain activity, we calculated the average network distance and spatial distance between different areas of the brain network. The optic and visual_projection area are related to vision and ALLN, ALIN, ALON, ALPN, LHCENT and LHLN are related to olfactory. The results are presented in Tab. 2.
From Tab. 2, it is evident that neurons which are closer in spatial distance are not necessarily easier to activate. For example, all areas and visual input neurons have similar average spatial distances, but the quasi-real activation patterns can only be observed in specific neural types (optic, visual_projection). In contrast, the average network distance between input neurons and neurons should be activated is closer than irrelevant neurons. For instance, the distance between visual input neurons and optic, visual_projection are 5.7018 and 5.7432, respectively. And the distance between visual input neurons and ALLN, ALIN, ALON, ALPN, LHCENT, LHLN which are irrelevant with vision are 6.6442, 6.2350, 6.1506, 6.5380, 6.3094 and 6.8206, respectively. The network distance of neurons related to visual input neurons is generally closer than the distances of unrelated neurons. Besides, The average network distance of neurons related to olfactory input neurons is approximately twice as close as the distance of unrelated neurons. These results indicates that compared to spatial distance, the average network distance between input neurons and their corresponding activation areas is closer, meaning that compared to spatial structure, network structure is more important, determining these quasi-real activation patterns of the brain network.
This results reflect the importance of network structure in neural networks, that is, the ability to reproduce certain functions of the brain by considering the network structure only.
Table 2: Spatial and network distance between input and main areas. The unit of spatial distance is nm and the unit of network distance is the number of edges in the shortest path.
distance type Neuron type Area Name optic visual_projection ALLN ALIN
Spatial Distance Visual Input Neurons $365012$ $354725$ $351919$ $348901$
Olfactory Input Neurons $302690$ $251144$ $110827$ $173345$
ALON ALPN LHCENT LHLN
Visual Input Neurons $346925$ $348517$ $362825$ $366013$
Olfactory Input Neurons $123179$ $115089$ $212515$ $246719$
optic visual_projection ALLN ALIN
Network Distance Visual Input Neurons $5.7018$ $5.7432$ $6.6442$ $6.2350$
Olfactory Input Neurons $5.7967$ $5.1706$ $3.4458$ $3.3845$
ALON ALPN LHCENT LHLN
Visual Input Neurons $6.1506$ $6.5380$ $6.3094$ $6.8206$
Olfactory Input Neurons $3.6312$ $3.6820$ $3.9557$ $3.9758$
3 Discussion
This study, through the construction of a large-scale network communication model based on the Drosophila connectome, reveals the critical role of neural network structure in generating brain-like activation patterns. Our findings suggest that models based on comprehensive neuron network data can approach the brain’s quasi-real activation patterns even with simple artificial rules. This paves the way for understanding biological intelligence, indicating that the structure of biological neural systems may offer valuable reference for designing efficient, fast-adapting intelligent systems.
Our research emphasizes the importance of network structure, compared to spatial structure or neuronal activation mechanisms, in determining brain activity patterns. This finding has significant implications for artificial neural network research, suggesting that future AI designs may need to more closely mimic real brain network structures.
Furthermore, the large spatial network visualization software we developed provides a new tool for neural network research, facilitating intuitive observation and analysis of complex neural network dynamics.
Future research directions may include optimizing propagation rules and automatically adjusting link weights to achieve complete simulation of brain behavior. These advancements will pave the way for achieving genuine artificial intelligence while deepening our understanding of the nature of biological intelligence.
4 Acknowledgement
We thank M.Murthy et al. for their invaluable help in sharing the Drosophila connectome data. This work was supported by the National Natural Science Foundation of China (72025405, 72088101), the National Social Science Foundation of China (22ZDA102), the Hunan Science and Technology Plan Project (2020TP1013, 2020JJ4673, 2023JJ40685), the Innovation Team Project of Colleges in Guangdong Province (2020KCXTD040), and the Postgraduate Scientific Research Innovation Project of Hunan Province (XJCX2023095). The authors declare that they have no conflict of interest.
APPENDIX
Appendix A Related Works
Various models have been developed to mimic or understand aspects of brain function, ranging from the cellular level to the whole brain. In this section, we review the models that aim to simulate the brain’s behavior. The review includes neural dynamic model that describes how action potentials in neurons are initiated and propagated. In addition, we introduce the large scale brain simulating methods. At last, we introduce the network communication models which refers to a computational or theoretical framework designed to understand and simulate how different parts of the brain communicate with each other.
A.1 Neuron Dynamic Model
Neuronal dynamic model is a mathematical framework designed to understand and simulate the electrical activity and dynamic behavior of neurons. These models aim to capture the fundamental mechanisms by which neurons process and transmit electrical signals through their membranes, often by using differential equations and computational techniques. In past years, neuroscientists have aimed to describing the behavior of neurons by dynamic model. One of the most famous works is the Hodgkin-Huxley (HH) model[18]. It includes a set of differential equations that describe the changes in membrane potential and the dynamics of ion channels. The key variables in the model are the membrane potential $V$, and the gating variables $m$, $n$, $h$ that represent the probability of ion channels being open.
The currents are given by:
$I_{Na}$ $=\bar{g}_{Na}m^{3}h(V-E_{Na}),$ (1)
$I_{K}$ $=\bar{g}_{K}n^{4}(V-E_{K}),$ (2)
$I_{L}$ $=\bar{g}_{L}(V-E_{L}),$ (3)
where $\bar{g}_{Na}$, $\bar{g}_{K}$ and $\bar{g}_{L}$ are the maximum conductances for sodium, potassium, and leak channels. $E_{Na}$, $E_{K}$, and $E_{L}$ are the reversal potentials for sodium, potassium, and leak channels.
The gating variables obey first-order kinetics and are described by the following differential equations:
$\frac{dm}{dt}$ $=\alpha_{m}(1-m)-\beta_{m}m,$ (4)
$\frac{dn}{dt}$ $=\alpha_{n}(1-n)-\beta_{n}n,$ (5)
$\frac{dh}{dt}$ $=\alpha_{h}(1-h)-\beta_{h}h,$ (6)
where $m$, $n$, and $h$ are the gating variables representing the activation and inactivation states of the ion channels. $\alpha_{m}$, $\alpha_{n}$, $\alpha_{h}$, $\beta_{m}$, $\beta_{n}$, and $\beta_{h}$ are voltage-dependent rate constants for the gating variables. These rate constants are defined by empirical equations based on experimental data. HH model accurately describes the electrophysiological processes of neuronal activity through mathematical equations, allowing scientists to understand the fundamental mechanisms of a group of neurons’ behavior at the molecular and cellular levels.
Another classic theory is cable theory[19]. Cable theory is a mathematical model used to understand how electrical signals propagate along neurons, especially in their dendrites and axons. This theory applies the principles of electrical circuits to describe how the passive properties of the neuron influence signal transmission. It’s particularly useful in explaining the conduction of graded potentials, such as synaptic potentials and electrotonic (passive) spread of membrane potential changes.
The core equation of cable theory is derived from the cable equation, which is a form of the second-order linear partial differential equation. It can be expressed as:
$\frac{\partial^{2}V}{\partial x^{2}}=\frac{1}{r_{m}}\left(\frac{\partial V}{\partial t}+\frac{V}{r_{c}}\right),$ (7)
where $V$ is the membrane potential, $x$ is the distance along the cable, $t$ is time, $r_{m}$ is the membrane resistance per unit length, $r_{c}$ is the cytoplasmic (axial) resistance per unit length. This equation describes how the voltage $V$ changes along the length of the neuron $x$ and time $t$.
Cable theory has several important applications in neuroscience, including understanding signal decay, neural modeling and synaptic integration. Despite its usefulness, cable theory has limitations. First, real neurons have complex geometries. Cable theory don’t describe well of these complex geometries. This can lead to inaccuracies in predicting real neuronal behavior. Besides, cable theory primarily addresses passive electrical properties and does not account for active properties like action potentials, which are crucial for neural communication.
Apart from that, researchers proposed many variants of neuron dynamic models[20][21]. These models serve to elucidate the physiological activity of neurons and contribute to the description of neuron behavior.
However, these models has some limitations: these equations often contain a multitude of parameters. In instances where measurement techniques fall short of quantifying each parameter individually, estimation becomes a necessity.
Spiking Neural Networks (SNNs)[22] is designed to optimize the parameters in neural dynamic models. Unlike their traditional counterparts, which process information through continuous values, SNNs utilize discrete events known as spikes. These spikes, brief in duration, mimic the communication method of neurons within the brain, facilitating the transmission of information.
One of the simplest and most commonly used models to describe neuron dynamics in SNNs[22] is the Leaky Integrate-and-Fire (LIF) model[23]. The LIF model captures the essential characteristics of spike generation in neurons. The membrane potential $V(t)$ of a neuron at time $t$ is described by the following differential equation:
$\tau_{m}\frac{dV(t)}{dt}=-[V(t)-V_{rest}]+RI(t),$ (8)
where $\tau_{m}$ is the membrane time constant, which determines how fast the membrane potential decays towards the resting potential $V_{rest}$ in the absence of input. $V_{rest}$ is the resting membrane potential. $R$ is the membrane resistance. $I(t)$ is the input current at time $t$. When $V(t)$ reaches a certain threshold $V_{threshold}$, the neuron fires a spike, and $V(t)$ is reset to $V_{rest}$ or another predefined value. After firing, the neuron may enter a refractory period during which it cannot spike, regardless of the input.
SNNs have enabled researchers to simulate networks comprising millions of neurons, tackling a wide array of tasks ranging from image classification and natural language processing to controlling robotic arms[24][25][26][27].
Despite their advances, SNNs face significant challenges. One primary issue is that the SNN’s structure often draws inspiration from functional brain networks rather than accurately replicating the intricate neural connections observed at the microscale. In terms of the connection method of artificial neurons, SNNs usually use a structure similar to that of ANNs. Consequently, these models resort to architectures akin to those found in artificial neural networks, hindering their ability to authentically emulate brain functions. Additionally, SNNs are marked by their low computational efficiency, with a single forward propagation frequently requiring several seconds to complete. These limitations underscore the difficulties in considering SNNs as the optimal approach for brain simulation in their current state.
As for the large scale brain simulation model aims to simulate the whole brain’s behavior, Eugene et al.[28] presents the first large scale brain simulation model based on experimental measures in several mammalian species. The model is founded upon the global thalamocortical anatomy derived from diffusion tensor imaging of the human brain. It incorporates multiple thalamic nuclei and a six-layered cortical microcircuitry, which is grounded in three-dimensional reconstruction of single neurons from the cat visual cortex. It boasts nearly half a billion synapses, incorporating accurate receptor kinetics, short-term plasticity, and dendritic spike-timing-dependent synaptic plasticity (dendritic STDP). The model can demonstrate behavioral patterns that mimic normal brain activity. This model is a comprehensive attempt to simulate the whole neural system, which is a critical component of the brain involved in sensory processing and consciousness.
Eliasmith et al.[29] introduce a more comprehensive brain simulating model, which consists of 2,500,000 SNN neurons. This model receives only sequences of visual images and produces all its responses through a simulated arm. Despite its simplifications of real structure, the model encapsulates numerous elements of neuroanatomy, neurophysiology, and behavioral psychology, as evidenced by its performance across eight distinct tasks.
As for the standardized platform, Zeng et al. [30] presents BrainCog, a cognitive intelligence engine inspired by the brain. This platform, based on SNN, offers critical infrastructure for the creation of brain-inspired artificial intelligence and simulation of the brain. BrainCog encompasses a variety of biological neurons, encoding methods, learning principles, brain regions, and a combination of hardware and software as its core elements.
However, large scale brain simulations require the simulation of millions or even billions of neurons, posing extremely high demands on computational resources. Even the most powerful supercomputers might need weeks or even months to complete a single simulation. Additionally, as the number of neuron connections and synaptic weights increases, data storage and processing requirements also significantly rise. Besides, although large scale brain simulations aim to simulate neurons and their connections, they often fail to capture all the biological details present in the brain. Factors like neuron morphology, molecular-level dynamics, and changes in gene expression are complex biological phenomena that are difficult to fully integrate into the models. Thus, the simulation results might not fully reflect the real behavior of the brain. At last, building a large-scale brain simulation model requires a vast amount of precise biological data, such as the connectivity between neurons, synaptic strengths, types, and concentrations of neurotransmitters. However, acquiring these data is often very challenging, involving complex and expensive experimental techniques, and it’s difficult to measure comprehensively across the entire brain.
As for the collective action of neuron populations, despite the success of neural system mean field models[31] in areas such as modeling epileptic seizures[32], there is still a lack of models that precisely describe the large-scale activity of whole brain.
A.2 Brain Network Communication Model
The network communication model is a conceptual framework which describes how data is transmitted from one node to another over a network. It breaks down the complex process of communication into simpler, more manageable layers or components. The most widely recognized network communication models are the Open Systems Interconnection (OSI) model and the Internet Protocol Suite, commonly known as the TCP/IP model. Nowadays, scientists apply network communication models to neuron research, which are mostly used to understand how brain networks operate[33, 34, 35].
Early studies have indicated that, similar to social and electrical networks[36], brain networks also exhibit properties such as power-law distributions[37] and the rich-club phenomenon[38]. These results indicate that brain networks contain a wealth of network statistical characteristics.
Researchers designed many brain network models to describe the processes of message propagation in brain. One category of brain network communication models is the routing protocol model. The most famous model of this category is shortest path model[39], which assumes that signals between a pair of nodes travel along the shortest path between them. In the shortest path model, the shortest path length from the starting node $s$ to each node $v$, is initialized as $d(s)=0$ and $d(v)=\infty$; then, process each node in increasing order of $d(v)$, for each node, update the $d(u)$ value of its neighbors $u$. The update way can be represented as follows:
$d(u)=\min(d(u),d(v)+w(v,u)),$ (9)
where $w(v,u)$ is the weight of edge $e_{uv}$.
Information is propagated entirely through the shortest paths in these models, which means in terms of propagation cost, signal transmission is optimal. However, the key disadvantage of the model is these methods require each neuron must have knowledge about the entire network structure[40], which doesn’t correspond to the reality: every neuron only has the message of its neighbor neurons.
Another well-known model is navigation model. It suggests that neurons always send information to the neuron that can be reached with the least communication cost[41]. These models achieve the lowest communication cost. A significant achievement is that some algorithms can identify paths with 70%-100% of optimal signaling efficiency in the functional connectomes of different organisms[42]. However, the main disadvantage of navigation model is it supposes that neurons know the distance between their neighbors and a desired target region. Recently their still has a lack of evidences that reveal how this knowledge would be given to neurons in brain networks.
Another well-known category of models is the diffusion model. The diffusion model posits that signals propagate through random walk dynamics. This process of propagation does not require individual neurons to possess any prior knowledge beyond their immediate neighboring neurons. Among the most classic of these is the random walk model[43]. In the random walk model, signals are transmitted to randomly selected neighboring nodes with a probability proportional to the connection weights until they reach the target node. For unbiased random walk, the function of random walk model is as follows.
$P_{ij}=\frac{1}{k_{i}},$ (10)
where $P_{ij}$ represent the probability of node $i$ choosing node $j$.$k_{i}$is the degree of node $i$.
As for biased random walk, for example, if the transition probability depends on the edge weights or certain properties of the nodes, then the transition probability can be adjusted to
$P_{ij}=\frac{w_{ij}}{\sum_{l\in N(i)}w_{il}},$ (11)
where $w_{ij}$ is the weight of node $i$ to node $j$,$N(i)$is the set of node $i$.
The diffusion process does not require individual neurons to have knowledge about the connectivity of connectome beyond their immediate neighbors. However, studies have shown that the efficiency of signal transmission through random walks is relatively low[44]. So it can’t describe the rapid communication process in the brain.
An effective category is threshold model[45]. The threshold model was originally proposed to characterize collective behavior. In the threshold model, a node’s behavior depends on the behavior of its neighboring nodes. Suppose that within a group, nodes can hold one of two states regarding a certain behavior: an active state and a quiescent state. For a node, it will only transition from the quiescent state to the active state if the stimulus given by its neighboring nodes in the active state surpasses a certain threshold; otherwise, it remains in the quiescent state. The condition for node $i$ to adopt the behavior at time $t$+1 can be formulated as follows.
$a_{i}(t+1)=1\quad\text{if}\quad\frac{1}{k_{i}}\sum_{j\in N_{i}}a_{j}(t)\geq\theta_{i},$ (12)
where $N_{i}$ is the set of neighbors of node $i$, $a_{i}(t)$ is the adoption state of node $i$ at time $t$. $\theta_{i}$ is the threshold of node $i$. $k_{i}$ is the degree of node $i$. The threshold model has stronger biological interpretability because it approximates the mechanism of neuronal firing. Consequently, the threshold model has been extensively applied in artificial neural networks. However, how to set the propagation thresholds for different neurons, and how to design the dynamics model between neurons remains a challenge.
Appendix B Data and Methods
B.1 Drosophila Connectome Data Set
We employ the recently released largest Drosophila brain connectome[17], offering a thorough mapping of the entire neural network within a Drosophila female brain. This data set encompasses over 120,000 neurons and 30,000,000 synapses, including detailed information on synaptic connections between neurons. Notably, it provides high-precision 3D coordinates, accurate to the nanometer level, for both cell bodies and synapses. Additionally, this data set offers extensive labeling for various types of neurons. For a more in-depth understanding, please refer to [17][46] for detailed information. In this section, we use nodes to represent neurons and connections to represent the synapses between neurons.
Table 3 presents the basic properties of the data set. Note that we combine multiple edges between two nodes to calculate network statistic properties. The average degree before and after combining multiple edges is 501.6 and 37.01, respectively, indicating it is prevalent that the pairs of neurons are connected by multiple synapses.
Table 3: Statistical properties of Drosophila Brain Connectcome Graph
Number of Nodes Number of Edges Average Degree
$131459$ $32970606$ $37.01$
Average Degree (before combination) Clustering Coefficient Eigenvector Centrality
$501.6$ $0.1527$ $0.0008817$
Network Diameter Assortativity Coefficient
$11$ $0.0008817$
B.2 Model
In this section, we introduce the large scale network communication model which consumes less computational resource and can efficiently utilize the network structure. To solve the problem that the link weights between neurons can’t be measured, we propose two hypotheses to estimate the link weights. In addition, we replace the node model in the network communication model with an improved LIF neuron model for comparison.
B.2.1 Symbol Systems
The original network presentation of neural systems often uses nodes to represent the neurons and using edges to represent the links (chemical synapses, electrical synapses, etc) between neurons[4]. However, this approach has a number of limitations: first, there are usually more than one link (synapse) between two neurons[49] and the links are spatially and temporally asymmetric. To illustrate, two neurons shown in Fig.S1 have multiple synapses between them, and the dynamical properties of these different synapses may vary. Additionally, due to their spatial locations, the signal transmission delay and the position of action on the downstream neurons are also different, making it an oversimplification if the synapses are merged into one edge.
Figure S1: An example of complex connection
For this reason, we use the dual method[50] on Drosophila connectome data set to model the multi-link between neurons.
In a dual graph, the nodes and edges of the original graph are interchanged, resulting in a new graph. In simple terms, a dual graph is obtained by replacing the nodes of the original graph with edges, and the edges with nodes. Fig. S2 is an example of an original graph and a dual graph. After being processed, the nodes in the dualed graph represent the synapses and links represent the cell body. Then we can use the communication model to describe the message propagation process.
(a) original graph
(b) dual graph
Figure S2: An example of original graph and dual graph
In order to standardize the symbol system, we use a edge (multi-edge, could include more than 2 nodes) to represent the neuronal cell body and use an node to represent a physical link between neurons (synapse). We regard the process of messages interactions as the nodes’ interactions via edges.
Let $N=\{n_{1},...,n_{i},...\}$ represents the set of all nodes and $E=\{e_{1},...,e_{j}...\}$ represents the sets of all edges. An obvious relationship is formalized as the following manner.
$e_{k}\cap e_{q}=\{n_{i}|(n_{i}\in e_{k})\wedge(n_{i}\in e_{q})\}.$ (13)
Note that the union of each element in $N$ and $E$ is the whole graph. The graph $G$’s definition is given as follows.
$G=\bigcup_{j}e_{j}=\bigcup_{i}\{n_{i}\}.$ (14)
Let the label of any neuron be $j$. Because neural link is directed, let ${e_{j}}^{+}$ ( ${e_{j}}^{-}$) represent the set of elements which output (input) the signals, note that ${e_{j}}$ = ${e_{j}}^{+}\cup{e_{j}}^{-}$.
Each neuron and synapse has an activate state (state = 1) or resting state (state = 0). When a neuron is at the activate state, it fires to other neurons, otherwise it keeps rest. So set the statement of the element $i$ as $X_{i}$. $X_{i}=1$ if the element $n_{i}$ is active, otherwise $X_{i}=0$. Then let $Y_{j}$ represent the statement of set $e_{j}$. $Y_{j}=1$ if the set $e_{j}$ is active, otherwise $Y_{j}=0$.
B.2.2 Large Scale Network Communication Model
By modeling the Drosophila connectome as graph model, we introduce the large scale network communication model on this graph. We use a linear threshold activation function on each neuron. Thus, general transmission function is defined in the following manner,
$Y_{j}=\sigma(\frac{1}{\beta_{j}}\sum_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j})X_{k}),$ (15)
where $\omega(n_{k},e_{i},e_{j})$ represents the link weight between set $i$ and set $j$ (specifically, two neurons) through element $n_{k}$. Here we consider that the link weight between two neurons is not only related to the neurons themselves, but also to the synapses between them. $\sigma(.)$ is the activate function. $\beta_{j}$ is the normalization factor.
In this paper we use threshold activation function and sigmoid function, which are two .
The threshold activation function is a simple, binary decision function used in artificial neurons. It outputs a value of 1 if the input exceeds a certain threshold, and 0 otherwise. This function is also known as the Heaviside step function.
$\sigma(x)=\left\{\begin{aligned} 1\quad x\geq threshold\\ 0\quad x<threshold\\ \end{aligned}\right.$ (16)
The sigmoid activation function is a smooth, S-shaped function that maps input values to an output range between 0 and 1. It is commonly used in neural networks to introduce non-linearity and to model probability-like outputs in binary classification problems. The sigmoid function is defined as the logistic function.
$\sigma(x)=\frac{1}{1+e^{-x}}$ (17)
The statement of an neuron will influence the statement of all output synapses. The equation between neuron $j$ and output synapse $i$ is as the following manner.
$X_{i}=Y_{j}\ \ \ \ if\ n_{i}\in e^{+}_{j}.$ (18)
The link weights between neurons determine the extent of influence they have on each other. Larger link weights mean a tighter ”link” between two neurons. In everyday life, the brain continuously changes the link weights between neurons in response to inputs from the external world, allowing intelligence to emerge from the micro to the macro. Therefore, if researchers want to simulate the process of information propagation in the brain on large scale in a computer, determining the link weights between neurons is essential. However, existing experimental methods are not well-suited for directly measuring the weights between neurons in a whole brain, so adopting some alternative approach is necessary. Below, we introduce a hypothesis that use different methods to estimate the connection weights between neurons.
As previously mentioned, there exists a large number of synapses between two neurons. However, in practical applications, it is difficult to set the parameters for these edges individually. Consequently, a simple assumption naturally arises: assuming that the source neuron has the same influence on the target neuron through every synapse. We propose an assumption that every synapse is equivalent, which is widely used in neuron graph statistics analysis[48]. That is to say,
$\omega(n_{k},e_{i},e_{j})=1.$ (19)
Additionally, We denote the sum of all link weights between neurons $i$ and $j$ as $W_{ij}$:
$W_{ij}=\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j}).$ (20)
And we define $\beta_{j}$ as the maximum weight among all $W_{ij}$ related to neuron $j$. That is to say,
$\beta_{j}={{Max}(\{W_{ij}|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\}).$ (21)
Where $Max(S)$ represents the Maximal element in set $S$.
Then the normalized equation is as follows.
$Y_{j}=\sigma(\frac{1}{\beta_{j}}\sum_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\omega(n_{k},e_{i},e_{j})X_{k}).$ (22)
B.2.3 Neuron Dynamic Model
To better reflect real-world conditions and demonstrate that the superiority of this network architecture does not solely rely on the design of the activation function, we also employed a variant of the LIF model[23] as the basic neuron model.
The membrane potential change of a neuron’s voltage can be described by the following differential equation.
$V_{j}(t+dt)=V_{j}(t)+\left(\frac{dt}{\tau}\right)\left(R\cdot I^{syn}_{j}-(V_{j}(t)-V_{\text{reset}})\right)$ (23)
Where $V_{j}$ represents the membrane potential of $j^{th}$ neuron, $R$ represents the resistance constant and $I^{syn}_{j}$ represents the total input current of neuron $j$. $\tau$ represents the time constant.
The input current $I^{\text{syn}}_{j}$ can’t be calculated directly. To estimate the value of $I^{\text{syn}}$, we make the following assumptions based on the following supposes:
1. The magnitude of the current on the synapse should be proportional to the voltage difference between the connected neurons.
2. The current on each synapse should be proportional to the connection weight between the neurons.
3. A current can only be emitted if the output neuron is in an active state.
Considering the supposes above, the equation of input current $I^{\text{syn}}_{j}$ is designed as follows:
$I^{\text{syn}}_{j}=\frac{1}{\alpha\beta_{j}}\sum_{{}_{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset}}\sum_{k|n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\left(\omega(n_{k},e_{i},e_{j})\times(V_{i}-V_{j})\times Y_{i}\right),$ (24)
where $V_{i}-V_{j}$ represents the voltage difference between neuron $i$ and $j$. $Y_{i}$ represents the state of output neuron $i$. $\beta_{j}$ is defined according to equation (21). $\alpha$ is the proportion const.
According to the LIF model, when a neuron’s voltage reaches the threshold value, the neuron will fire and then quickly resets its voltage. Based on this, the mechanism for updating the neuron voltage is designed as follows.
$\text{if}\ V_{j}\geq V_{th}&:\ Y_{j}=1\ \text{and}\ V_{j}=V_{\text{rest}}\\ \text{else}&:\ Y_{j}=0\end{aligned},$ (25)
where $V_{th}$ represents the fire voltage and $V_{\text{rest}}$ represents the rest voltage, respectively.
Other propagation mechanism is consistent with network model introduced in section B.2.2.
B.3 Evaluation Criteria
In previous study, researchers used classic indicators such as signaling cost and activation time to evaluate the quality of the model[16][51][52]. However, these indicators can’t describe the model accords with the true communication process or not, even if the activation time is short and the signaling cost is minimum, etc. To better assess whether a model can reflect the real situation of brain communication, we use the percentage of activated neurons in regions that related to (should work under) a certain stimulus. To illustrate, if the brain gets a visual stimulus, the region related to vision should work while the region not related to vision should remain a resting state. Then we record the percentage of neurons activated. If regions related to the certain function have a significant higher activation percentage than the average activation percentage of the whole brain, we call it a quasi-real activation pattern.
We define the consistency indicator $C_{r}$ in region $r$ which has a certain function (visual, olfactory, etc) is defined in the following manner.
$C_{r}=N_{r}^{act}/N_{r},$ (26)
where $N_{r}^{act}$ represents the number of neurons activated in area $r$ at the final state, $N_{r}$ represent the number of neurons in area $r$ and $N$ represent the total number of neurons.
Appendix C Introduction of Related Neurons
Optic neurons, also known as retinal ganglion cells, are the first neurons in the visual pathway. They receive signals from the photoreceptor cells in the retina and transmit this information via the optic nerve to the brain. These neurons play a crucial role in carrying visual stimuli from the eye to the brain, where further processing and interpretation of the visual information occur.
Visual projection neurons are a type of neuron found in the visual system that transmit visual information from one region of the brain to another. They play a crucial role in processing and relaying visual signals from the retina to various visual centers in the brain, including the thalamus, hypothalamus, and primary visual cortex (V1), among others.
Antennal Lobe Input Neurons (ALINs) are primarily olfactory sensory neurons (OSNs) that carry olfactory information from the sensory structures, such as the antennae and maxillary palps, directly to the antennal lobe. Each OSN expresses a specific odorant receptor and responds to particular chemical cues. The axons of these neurons synapse with projection neurons and local neurons within the glomeruli of the antennal lobe.
Antennal Lobe Projection Neurons (ALPNs) transmit the olfactory information in the antennal lobe. These neurons receive input from the olfactory sensory neurons within the glomeruli and project their axons to other brain areas, such as the mushroom body and the lateral horn, where further processing and integration of olfactory information occur.
Antennal Lobe Output Neurons (ALONs) refer to the projection neurons that serve as the primary output pathways of the antennal lobe, sending processed olfactory information to higher brain regions.
Antennal Lobe Local Neurons (ALLNs) are inter neurons that are confined to the antennal lobe. They typically have processes (dendrites and axons) that branch extensively within the antennal lobe and form connections with multiple glomeruli. These neurons play a critical role in modulating and refining the olfactory information by providing inhibitory or excitatory input to projection neurons and other local neurons, thus shaping the olfactory responses and contributing to odor discrimination and perception.
For further information, please read [46].
Appendix D Bilateral Response under Unilateral Stimulus
In real life, the left and right hemispheres cooperate to perform complex tasks. We are interested in whether network models can simulate this cooperative effect to some extent. So we administer a unilateral stimulus to the input neurons and monitor the varying activation ratios across both hemispheres. More precisely, we independently simulate visual input neurons in the left and right hemispheres and record the rate of activated neurons in each hemisphere over time. Fig.S3 illustrates the curve depicting how the response ratio evolves with each time step. The horizontal line represents the proportion of activated neurons in all hemisphere. It is noteworthy that when simulating one hemisphere alone, visual processing neurons in both hemispheres are activated. At time step 5, the unstimulated hemisphere begins to have a activate pattern, which corresponds to the hypothesis of shallow connection[53]. Furthermore, our observations reveal a concurrent effect: the activation pattern in one hemisphere mirrors that in the unstimulated hemisphere.
These findings are reflective of reality. Although the left and right hemispheres of the brain possess certain functional specializations, they also demonstrate a synergistic effect. Typically, they collaborate to fulfill specific functions[56]. For instance, during complex cognitive tasks, both hemispheres coordinate their efforts to contribute to task completion. Even when simulating input to neurons on just one side of the brain, our model reveals similar synergistic effects, indirectly attesting to the model’s effectiveness. Besides, the mirror effect align with real situation, underscoring the brain’s resilience as a system. In instances of disease or physical damage leading to impairment in one hemisphere, this interconnected structure supports the continuation of critical functions, such as vision.
(a) Left stimulus
(b) Right stimulus
Figure S3: Experiment of activation rate under single stimulus. We set threshold $\sigma_{0}=0.8$.
Appendix E Impact of Perturbation Rate on Activation Areas
The main text demonstrates that with a 10% perturbation in the network, the original realistic activation patterns vanish regardless of the model used. In this section, we discuss the impact of different reconnection rates on activation patterns using the degree-preserving reconnection method.
The experiments presented in this section involve calculating patterns induced by two types of stimuli across three models under network perturbations of 1%, 3%, 5%, 10%, and 20% reconnect rates. We normalized each data point using the difference between the activation ratio of region $r$ and the average activation ratio $\bar{C}$ of the entire brain, that is to say, $\Delta C=C_{r}-\bar{C}$.
The experimental results are shown in Fig.S4, where 0 represents the original, unperturbed experimental results. It can be observed that at the 0 point, there is a significant difference in the relative activation ratios between visual and olfactory areas under specific stimulus. However, this significant realistic activation patterns which we discussed in the main text disappear with a 1% network perturbation, replaced by a roughly equal activation ratio across all regions. This indicates that even a 1% change in network structure can lead to a rapid loss of network functionality specificity. Additionally, it is noted that small-scale network perturbations (1%-5%) have a lesser impact on olfactory stimulus compared to visual stimulus. This is because the number of neurons associated with olfactory functions is significantly fewer than those related to visual functions, making it less likely for small-scale network structure changes to affect these neurons, thereby leading to a smaller impact on functionality.
At the same time, various horizontal comparisons indicate that while there are no significant differences among different models in forming activation patterns, more complex neuron models have a stronger ability to maintain their original activation patterns under network disturbances. This is because the signals generated by complex neuron models have greater redundancy and functional diversity. For instance, the LIF (Leaky Integrate-and-Fire) model has a voltage decay mechanism, where the voltage gradually decays in the absence of a signal. Even if some input signals are erroneous, as long as the proportion of errors is not very large, the decay mechanism can filter out the impact of these erroneous signals.
These results indicate that the formation of realistic activation patterns is highly sensitive to changes in network structure, with even minor perturbations capable of disrupting this network configuration. This indirectly substantiates that the authentic network structure is crucial for the formation of genuine activation patterns. Additionally, it highlights that complex functions (such as vision) are more susceptible to network disturbances.
(a) threshold model, visual stimulus
(b) sigmoid model, visual stimulus
(c) LIF model, visual stimulus
(d) threshold model, olfactory stimulus
(e) sigmoid model, olfactory stimulus
(f) LIF model, olfactory stimulus
Figure S4: Relationships of activation rate with reconnect rate. We conduct 5 repeated experiments and take the average and standard deviation.
Appendix F Large Spatial Network Visualization
The visualization software of experiment observation on large spatial networks is important for researchers[57][58]. However, there is still a lack of visualization system that can observe information propagation on large spatial networks in real-time, which poses difficulties in experiment. To observe directly the communication progress on this super large Drosophila connectome, we develop a large spatial network visualization software. This software will be made available to other researchers with the author’s permission.
The drawing part mainly employs the Canvas provided by HTML5 and we use the Three.js based on WebGL technology to achieve spatial network visualization effects, while the user interface display is constructed by Vue.js, providing users with an intuitive and interactive interface to display and analyze data.
F.1 Front End Technical Details
We use the Vue.js framework to build the user interface. Vue.js is a popular JavaScript framework that simplifies the development of web applications through componentization, enabling the creation of front-end interfaces with rich interactivity and good user experience.
The rendering of three-dimensional complex networks relies heavily on the Three.js library. Three.js is a powerful JavaScript library for creating and displaying 3D graphics. It offers a rich set of 3D objects, materials, lights, and animations, allowing for the easy creation of high-quality 3D scenes and objects.
To improve rendering performance, GPU acceleration technology is utilized. GPU acceleration takes advantage of the computational power of the Graphics Processing, significantly enhancing rendering speed and efficiency through parallel processing and stream processing modes. The GPU acceleration capabilities of Three.js, especially through the creation of point objects using Float32BufferAttribute, which contains their position information in three-dimensional space, greatly enhance the rendering speed and smoothness.
F.2 Back End Technical Details
We use the Spring Boot framework, an open-source framework based on Java for quickly building independent, production-grade Spring applications. With Spring Boot, backend services can be rapidly set up to provide RESTful API interfaces for frontend consumption. The backend also utilizes the PostgreSQL database to store network data.
Besides, We use PostgreSQL, which is a powerful open-source relational database management system that offers a rich set of data types and robust querying capabilities. To store 3D network data, we leverage PostgreSQL’s JSONB type, which allows storing JSON-formatted data within PostgreSQL and provides a series of functions and operators for querying and manipulating this data. With the JSONB type, information about nodes and edges in the network can be stored, enabling fast querying and updating operations.
F.3 Function Display
Fig.SS5 is an example of the 3D visualization system.
(a) Side view 1
(b) Side view 2
(c) Side view 3
(d) Top view 1
(e) Top view 2
(f) Top view 3
(g) Front view 1
(h) Front view 2
(i) Front view 3
Figure S5: Visualization of Spatial Drosophila Brain Network
All nodes’ colors can be adjusted by visible selection or input RGB code (Fig.SS6).
Figure S6: Display of color selection
Besides, from square to round, the shape of nodes can be adjusted even real neuron (Fig.SS7).
Figure S7: Display of color selection
It also can present different single areas (Fig.SS8).
Figure S8: Display of color selection